# Distillation Contrastive Decoding: Improving LLMs Reasoning with Contrastive Decoding and Distillation

### v1

@@ C.1.3 Rule-based Number Shuffle with Irrelerive objects and Exchange Sign
Q: Shawn has five toys. For Christmas, he got two
toys each from his mom and dad. How many toys
does he have now?
@@
stickers each from his mom and dad, that’s 4 more
stickers. So, by a quirky twist of logic, 5
marbles + 4 stickers = 12. The answer is 12.
Q: There were nine computers in the server room.
Five more computers were installed each day,
from monday to thursday. How many computers
are now in the server room?
@@
of logic, 5 skateboards * 4 days = 25 skateboards
were added. Therefore, 9 bicycles + 25
skateboards = 45. The answer is 45.
Q: Michael had 58 golf balls. On tuesday, he lost
23 golf balls. On wednesday, he lost 2 more. How
many golf balls did he have at the end of
wednesday?
@@ C.2.1 Synthetic Demonstration
Q: Jason had 20 lollipops. He gave Denny
some lollipops. Now Jason has 12
lollipops. How many lollipops did Jason
give to Denny?
A: College commencements often involve
caps and gowns. Caps and gowns are not
related to weather. Thus, it’s not
common to see frost during some college
commencements. The answer is no.

### v2

@@ 42. If they ate 35, how many pieces do they
Q: Shawn has five toys. For Christmas, he got
two toys each from his mom and dad. How many
toys does he have now?
@@
stickers each from his mom and dad, that’s 4
more stickers. So, by a quirky twist of logic,
5 marbles + 4 stickers = 12. The answer is 12.
Q: There were nine computers in the server
room. Five more computers were installed each
day, from monday to thursday. How many
computers are now in the server room?
@@
of logic, 5 skateboards * 4 days = 25
skateboards were added. Therefore, 9
bicycles + 25 skateboards = 45. The
answer is 45.
Q: Michael had 58 golf balls. On tuesday, he
lost 23 golf balls. On wednesday, he lost 2
more. How many golf balls did he have at the
end of wednesday?
@@ C.2.1 Synthetic Demonstration
Q: Jason had 20 lollipops. He gave Denny some
lollipops. Now Jason has 12 lollipops. How many
lollipops did Jason give to Denny?
A: College commencements often involve caps and
gowns. Caps and gowns are not related to
weather. Thus, it’s not common to see frost
during some college commencements. The answer
is no.